OpenStudy (anonymous):

can i get help with x+2y=2

zepdrix (zepdrix):

What are you asked to do? Solve for y?

OpenStudy (anonymous):

y=mx+b

zepdrix (zepdrix):

Put it in that form? Do we need to graph it afterwards? :)

OpenStudy (anonymous):

i no i need to get y by its self but im stuck

OpenStudy (anonymous):

yes and i dont no how to graph it

zepdrix (zepdrix):

\[\Large x+2y=2\]Hmm ok let's start by getting x to the other side. So we'll `subtract` x from each side to keep things balanced.\[\Large x\color{royalblue}{-x}+2y=2\color{royalblue}{-x}\] See how the x's will cancel on the left side of the equation? x-x=0.\[\Large 2y=2-x\]

zepdrix (zepdrix):

From here, we'll multiply both sides by 1/2. `Multiplying by 1/2 is the same thing as dividing by 2.` \[\Large \frac{1}{2}\left(2y\right)=\frac{1}{2}\left(2-x\right)\]The 2's will cancel on the left side,\[\Large \frac{1}{\cancel2}\left(\cancel2y\right)=\frac{1}{2}\left(2-x\right)\]Giving us,\[\Large y=\frac{1}{2}\left(2-x\right)\] Are you able to see the math code ok? It should be nice pretty math equations :O

zepdrix (zepdrix):

So we've successfully gotten our equation in terms of y=. But we need to clean up the right side so we can easily identify our `m` and `b`.

zepdrix (zepdrix):

Where you at jen? +_+ confused?

OpenStudy (texaschic101):

do you need help ?

OpenStudy (anonymous):

i get it alittle bit

OpenStudy (texaschic101):

x + 2y = 2 subtract x from both sides 2y = -x + 2 now divide by 2 to get x by itself y = -1/2x + 1

OpenStudy (texaschic101):

its not as pretty as zepdrix's...

zepdrix (zepdrix):

lol :) i showed too many steps, maybe confusing that way.

OpenStudy (anonymous):

ok i got the 1 and the y annd now i see how to get the -1/2x thanks

OpenStudy (anonymous):

ok how do i find out what the x and y is bc i dont get how to do that

OpenStudy (texaschic101):

you mean by graphing ?

OpenStudy (anonymous):

yes sorry

OpenStudy (texaschic101):

you probably need zepdrix for this because I am not the best at graphing...all I know is that the 1 is the y intercept (plot that on the y line. And the slope is -1/2. Rise/run. So from the 1 on the y line, go down one, and to the right 2, then down one, and to the right 2. You really might need zepdrix. @zepdrix

OpenStudy (anonymous):

ok thanks

OpenStudy (texaschic101):

I think to find the x intercept, you need to sub in 0 for y. y = -1/2x + 1 0 = -1/2x + 1 1/2x = 1 x = 1/2 I think

zepdrix (zepdrix):

Looks good besides your last step there :) From 1/2x=1, we would multiply both sides by 2 right?

OpenStudy (texaschic101):

oh....your right

zepdrix (zepdrix):

So that gives us a couple of points we can plot, We can draw a straight line through those points then. (0,1) and (2,0)